For normal incidence of light from a medium of refractive index 1.42 onto a medium of refractive index 1.64.calculate the refection and transmission coefficents.
1
Expert's answer
2020-08-31T12:47:20-0400

solution:-

given data

refractive index of medium 1 "(n_1)" =1.42

refractive index of medium 2"(n_2)=1.64"

the reflection coefficient can be written

"R=\frac{(n_2-n_1)^2}{(n_2+n_1)^2}"


"=\frac{(1.64-1.42)^2}{(1.64+1.42)^2}"


"=\frac{(0.22)^2}{(3.06)^2}"

"\fcolorbox{red}{yellow}{$R=0.0052$}"

therefore reflection coefficient is 0.0052

relation between reflection and transmission coefficient can be written as

"T+R=1"

"T=1-R"

by putting the value of R

"T=1-0.0052"

"\fcolorbox{red}{yellow}{$T=0.9948$}"

therefore transmission coefficient is 0.9948 .